# Model Compression in Practice: Lessons Learned from Practitioners Creating On-device Machine Learning Experiences

### v1

@@ 3. A Primer on Model Compression Techniques
To familiarize readers with model optimization in general, here we give a primer on common ML compression techniques. This background will help ground the study results and provide context for the remainder of the paper.
@@
Model compression is a class of techniques used in on-device ML that optimizes varying components of a model to minimize their required computational resources. While it is not important to know every detail of every technique, it is useful to understand of the variety of techniques, how they work at a high-level, and how they can be combined together. In fact, some of the biggest savings can come from effectively combining multiple techniques (Han et al., 2016). This overview contains a brief description of each technique, and includes illustrations to build visual intuition. We only attempt to cover techniques mentioned by practitioners in our study. Each technique below is truly a family of techniques, each with many nuanced variations. For an in-depth review of the technical descriptions of compression techniques, see the following surveys: (Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018).
@@ 3.2. Palettization (Weight Clustering)
Map the weights of a model to a discrete set of precomputed (or learned) values (Cho et al., 2022; Wu et al., 2018). Inspired by an artist’s palette, the idea is to map many similar values to one average or approximate value, then use those new values for computing inference. In this way, palettization is similar in spirit to algorithmic memoization, or more simply, a dictionary or a look-up table. Palettization can make a model smaller but does not make a model faster since it incurs look-up time.
3.3. Pruning (Sparsity or Network Sparsity)
Remove the least-important parts of a neural network to make it smaller (Hoefler et al., 2021). The motivation here is that modern ML networks are much more dense than is actually needed. Since networks can contain millions of parameters, removing some or many parameters may not impact the final accuracy. Pruning is a large family of techniques and strategies. For a detailed overview see (Hoefler et al., 2021).
@@ 3.3.1. Unstructured Pruning
A model may have neurons or weights that do not highly affect a model’s decision (Gholami et al., 2021). In unstructured pruning, the least important neurons or weights can be aggressively removed without affecting model accuracy (Hoefler et al., 2021). Unstructured pruning has been shown to shrink a model by 10–100x its original size (Hoefler et al., 2021). The major downside is that this leads to sparse matrix operations, which is a type of math that is slow on most hardware. So although pruned models are small, they may be nearly as slow as the full-sized model (Gholami et al., 2021).
@@ 3.3.2. Structured Pruning
Similar to unstructured pruning, the least important neurons or weights are identified, however, instead of removing (or zero-ing out) individual values, structured pruning removes entire structural elements, like channels or filters. Since structured pruning removes much bigger chunks of a model, less pruning will be possible without serious degradation in model accuracy (Gholami et al., 2021). However, by respecting the structure of the model, these pruned models maintain dense matrix operations, which make them fast on most hardware (Hoefler et al., 2021).
@@ 3.4. Distillation
Train a smaller model to mimic a larger model. Given a large (teacher) model that is already highly accurate at a task, the idea is to transfer the “knowledge,” or learned features, to a smaller (student) model by having the smaller model replicate the larger model’s output. Extending this further, given a newly trained student model, one can then apply other compression techniques to the student model for further optimization (Gholami et al., 2021; Polino et al., 2018). For a survey of distillation techniques see (Gou et al., 2021).
@@ 3.5. Efficient Neural Architectures
Some model architectures are specifically designed to be small and efficient. One of the most prominent examples in recent years is MobileNets (Howard et al., 2017; Sandler et al., 2018), with others like ShuffleNets (Zhang et al., 2018), or EfficientNets (Tan and Le, 2019). All of these examples are designed to improve the efficiency of convolutions, which are vital to fitting computer vision models on-device (Choudhary et al., 2020). Efficient architectures can be compressed with techniques like quantization or pruning, to create an even smaller model (Choudhary et al., 2020).
@@ 3.6. Dynamic Models
When thinking about ML applications, most people think of a single model with fixed inputs and outputs. Dynamic models differ in that they take into account the the differing prediction difficulty of each data instance. In other words, based on the individual data point to run inference on, a model may adapt its preprocessing steps, computational path, or model choice all together (Zhu et al., 2021). While many types of dynamic models exist, we only discuss a selection of techniques that are mentioned later on.
@@ 3.6.1. Early Exit Models
Allow for completing a prediction without the need to pass through the full model. Motivated by the fact that some data points are easier to predict than others, early exit models run intermediate feature representations through additional output layers to check if the prediction is sufficiently correct. For example, one way to determine whether or not to early exit or continue through a network is to check its prediction confidence: if a model is already confident about the current prediction, early exit, and if not, continue to the next prediction layer. At a high-level, these are models with checks throughout an architecture, such that inference can complete at any check before the final output.
@@ 3.6.2. Gated Models
Train a smaller, approximate model whose prediction decides whether or not to invoke a larger model. Similar to early exit models, some data points are easier to predict than others. A fast, smaller model gates a larger model, for example by again using the confidence of the initial prediction. At a high-level, these are systems of models that decide whether or not to spend extra compute if a prediction is still unknown.
@@ 4.1. Study Protocol
To capture emerging practices around efficient machine learning, we conducted semi-structured interviews (Boyce and Neale, 2006; Knott et al., 2022) with 30 ML practitioners at Apple to study how they approach model compression in their own projects. Our interview questions are outlined in Appendix A. We gave participants ample time to flexibility speak to their specific work and express other topics beyond our question set that they felt were important to doing good efficient machine learning work (Knott et al., 2022). The interviews took place between March and July 2022 with each conversation lasting from 45 minutes to 1 hour. For all interviews, one researcher lead the interview questions, while another took detailed notes. Where a participant approved, we also recorded conversations to refer back to during analysis. No compensation was given, since all participants were salaried employees. At the end of the study, we briefed participants and their teams on our results. Our study protocol was approved by an internal IRB.
@@ 4.2. Participants and Recruitment
As discussed in Section 2, in-depth interview studies with ML practitioners are uniquely suited to capture experts’ tacit knowledge for the purpose of generating new publicly-available guidance (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Our organization is one of the few places in the world that contains a concentration of efficient machine learning experts, so we first reached out to known experts. We then used a snowball sampling strategy to reach a broader network of people involved in efficient machine learning at Apple, including other compression experts, ML practitioners who own a production model but are not compression experts themselves, and tooling engineers who build frameworks for model compression. To balance different perspectives, we sought practitioners working on ML for different domains (e.g., vision, text, and sensor data) and different use cases. As the study evolved, we also chose new participants to help fill-in our biggest areas of uncertainty. Our participants, listed in Section 5.1, include ML research scientists, engineers, and managers spanning a wide breadth of application domains. A natural saturation occurred when participants began to repeat know-how we had already recorded and began to only suggest participants we had already included.
@@ 4.3. Qualitative Data Analysis
The first two authors conducted an iterative thematic analysis method to group common motivations, challenges, and best practices of model compression into categories (Gibbs, 2007). As we conducted more interviews, we continually discussed and updated these categories to reflect our new findings. Each participant’s data and transcripts were independently reviewed and manually coded by the first two authors at least twice using inductive coding (Thomas, 2003). In total, we spent 30+ hours interviewing participants and collected 23,500+ words of additional notes outside of the audio transcripts. The final categories were split into two main sections: Section 5 “Study Results” and Section 6 “Design Opportunities for Efficient Machine Learning Interface.” The composition of these two sections was formed by ordering major categories and hierarchically grouping within categories when relevant.
@@ 4.4. Methodological Limitations
While we found that learning from expert practitioners was tremendously valuable, any interview study has limits for generalizability. Interview studies suffer from smaller population samples than other methods, for example, a survey. We chose to hold in-depth exploratory conversations with each participant—which a survey cannot support as well (Knott et al., 2022). On the other hand, we did not directly observe participants conducting their efficient machine learning work, which limits the specificity of our findings. It was not possible to observe participant’s direct work or model artifacts due to the sensitivity of their work. Another concern is bias from a participant’s role in the domain (Boyce and Neale, 2006), as well as the power dynamics between interviewer and interviewees (Kvale, 2006). This study was conducted solely within one organization, therefore practitioners may hold organization-specific beliefs and practices (Schein, 1990). Despite our conscious efforts to recruit across ML application domains, we noticed a skew towards vision-based applications on images, video, and 3D scene data. This makes sense from an industry perspective, as computer vision applications typically rely on exceptionally large ML models. Lastly, these interviews were conducted prior to the public rise of LLMs in late 2022 (Movva et al., 2023). As discussed in subsection 2.7, the same overall approaches for efficient machine learning are currently being applied to LLMs and other foundation model modalities.
Taking these limitations together, we caution readers to not consider the advice we detail from participants universally generalizable. However it is with confidence that we share the rich pragmatic advice, guidance, and experience these 30 experts have to offer.
@@ 5.1. Participants and Emergent Personas
We summarize our participants in Section 5.1. Of the 30 people we interviewed, participants had an average of 7.1 (max 12) years experience with ML. Participants had 4.1 (max 10) years on average experience with efficient ML. Our participants had diverse breadth of expertise across domains and mapped roughly to four roles: 6 ML Managers, 4 Research Scientists, 13 ML Engineers, and 7 Software Engineers. For the purpose of understanding practitioners roles relative to efficient machine learning, we found three distinct personas emerged that best describe our participants:
@@
A full-page table divided into three sections, each describing one of the personas of the study participants. Each row represents a participant, and lists their years of experience, job title, and what ML application they focus on.
@@
[2pt/2pt] E1 10/12 ML Manager (✓) Efficient ML research
@@
[2pt/2pt] P1 6/10 ML Manager (✓) Efficient computer vision
@@
[2pt/2pt] T1 5/10 ML Manager (✓) Efficient ML tooling
@@
  • Compression Experts (E1–E13 in Section 5.1): Seasoned research scientists and experienced engineers that lead on-device ML initiatives, develop new compression techniques, and consult on machine learning optimization efforts. Given the amount of experience these people have, they often manage or lead teams. Example: An ML research scientist who has a PhD in model compression and is developing novel techniques for model optimization.
  • Machine Learning Practitioners (P1–P11 in Section 5.1): ML engineers and data scientists that build and deploy models on-device where certain optimization budgets must be met. These people use compression as a means to an end rather than solely studying efficient machine learning. Example: An ML engineer needs to get a model’s size under 1MB while maintaining high accuracy.
  • Tooling Engineers (T1–T6 in Section 5.1): Engineers and developers that focus on building frameworks, infrastructure, and tools for efficient machine learning. Example: A software engineer that builds and maintains an organization-wide tool to help others compute efficiency metrics.
Throughout the paper, we use representative quotes from participants to illustrate the main findings from the study.
5.2. The State of Efficient Machine Learning
Creating an efficient machine learning model is a “large design and constrained optimization problem” [E5], where practitioners have to weigh decisions between many model performance metrics, such as model storage size, inference latency, power usage, device temperature, and model behavioral metrics such accuracy, precision, recall, and others. To many readers, this characterization may sound like a job for an automated optimization algorithm. However, a crucial finding of our results is that automation is not yet possible, do to the degree of both ML and product design human expertise that goes into creating on-device applications. Experts emphasized there is no single solution, “silver bullet” [E12], “turnkey solution” [E5], or “golden recipe” [E7] for successfully building efficient machine learning models. Partially, this is due to the rapidly moving-target of new state-of-the-art model architectures and new ML applications that may require a custom approach.
@@ 5.2. The State of Efficient Machine Learning
A full-width table with two columns: user perceivable metrics on the left and negative user impact on the right. The user perceivable metrics column contains colored icons to indicated metrics like model storage size as a box, power usage as an empty battery, and device temperature as a flame.
@@
[2pt/2pt] All metrics: model tested during high device resource load Resources are shared. A user’s device may be running many things simultaneously, including multiple ML models. A single model that demands too many resources will slow everything down.
The expertise for creating efficient ML models is currently held by a relatively small subset of ML practitioners, where “knowledge is handed-down from the few who know how to do it successfully,” [E3]. These people are referred to as artisans and “based on earlier experience, know it’s possible” [P10] to create efficient models. Reducing a model’s size by half, or by an order of magnitude, takes additional time and considerations [E7]. “It’s like black magic or the dark arts, but you will get better with time,” [E7]. Readers may recognize these “dark arts” or “artisan” language from earlier days of contemporary ML (Patel et al., 2008; Hill et al., 2016) before any real widespread knowledge base was established. We hope these study results support growing such a knowledge-base for efficient machine learning, and it’s role within a holistic ML process:
[Efficient machine learning] means a lot of things to different people. Specific techniques on models weights help reduce size, but to get an efficient model comes from more careful design of the loss function, the system, which parts should and should not be modeled with ML.” — E5
To this end, we think there is rich opportunity to help share and inform practice. Our interview study results characterize how practitioners design on-device machine learning experiences (5.3), decide model budgets (5.4), invest in compression (5.5), and evaluate compressed models to balance accuracy versus performance and avoid compression artifacts (5.6). We conclude with distilled design recommendations for tooling that could address the challenges practitioners face and bring efficient machine learning practices mainstream (Section 6). Throughout the results we highlight summarized takeaway strategies in call-out boxes.
5.3. The Metrics of Efficient, On-Device Models
@@
As an umbrella term, efficiency encompasses many metrics. Common metrics that practitioners linked to user experience are summarized in Table 2. Some may be familiar with metrics commonly considered in ML workflows, such as latency and model storage size. Other metrics may come as more foreign to an ML engineer accustomed to deploying their model on a far-off server, e.g., device temperature. When a user is holding and carrying around a device all day, metrics like heat and battery life impact are crucial.

### v2

@@ 3. A Primer on Machine Learning Compression Techniques
To familiarize readers with model optimization, here we give a primer on common ML compression techniques. This background will help ground the study results and provide context for the remainder of the paper. Note we only cover common techniques mentioned by practitioners in our study—more exist.
@@
Model compression is a class of techniques used in on-device ML to reduce the computational resources a model consumes. While it is not important to know every detail of every technique, it is useful to understand the variety of techniques at a high-level, and how they can be combined for bigger savings (Han et al., 2016). This overview contains a brief description of each technique, and includes illustrations to build visual intuition. Note that each technique below is truly a family of techniques, each with many nuanced variations. For an in-depth review of the technical descriptions of compression techniques, see the following surveys: (Menghani, 2023; Deng et al., 2020; Choudhary et al., 2020; Cheng et al., 2018).
@@ 3.2. Palettization (Weight Clustering)
Map the weights of a model to a discrete set of precomputed (or learned) values (Cho et al., 2022; Wu et al., 2018). Inspired by an artist’s palette, the idea is to map many similar values to one average or approximate value, then use those new values for computing inference. In this way, palettization is similar in spirit to algorithmic memoization, a dictionary, or a look-up table. Palettization can make a model smaller but does not make a model faster since it incurs look-up time.
3.3. Pruning (Network Sparsity)
Remove the least-important parts of a neural network to make it smaller (Hoefler et al., 2021). The motivating idea is that modern neural networks are much more dense and overparameterized than is actually needed. Since networks can contain billions of parameters, removing some or many parameters may not impact the final accuracy. Pruning is a large family of techniques (Hoefler et al., 2021).
@@ 3.3.1. Unstructured Pruning
A model may have neurons or weights that do not much affect a model’s decision (Gholami et al., 2021). In unstructured pruning, the least important neurons or weights can be aggressively removed without affecting model accuracy (Hoefler et al., 2021). Unstructured pruning has been shown to shrink a model by 10–100x its original size (Hoefler et al., 2021). The major downside is that this leads to sparse matrix operations, which is a type of math that is slow on most hardware. So although pruned models are small, they may be nearly as slow as the full-sized model (Gholami et al., 2021).
@@ 3.3.2. Structured Pruning
Similar to unstructured pruning, the least important neurons or weights are identified, however, instead of removing (or zero-ing out) individual values, structured pruning removes entire structural elements, like channels or filters. Since structured pruning removes much bigger chunks of a model, less pruning will be possible without serious degradation in model accuracy (Gholami et al., 2021). However, by respecting the structure of the model, these pruned models maintain dense matrix operations, which keeps the model fast on most hardware (Hoefler et al., 2021).
@@ 3.4. Distillation
Train a smaller model to mimic a larger model. Given a larger (teacher) model that is already highly accurate, transfer its learned features to a smaller (student) model by having the smaller model replicate the larger model’s output. One can then apply other compression techniques to the student model for further optimization (Gholami et al., 2021; Polino et al., 2018). For a survey of distillation techniques, see (Gou et al., 2021).
@@ 3.5. Efficient Neural Architectures
Some model architectures are specifically designed to be small and efficient. Prominent example architectures include MobileNets (Howard et al., 2017; Sandler et al., 2018), ShuffleNets (Zhang et al., 2018), and EfficientNets (Tan and Le, 2019). All of these examples are designed to improve the efficiency of convolutions, which are vital to fitting computer vision models on-device (Choudhary et al., 2020). One can use other compression technqiues on efficient architectures to create even smaller models (Choudhary et al., 2020).
@@ 3.6. Dynamic Models
When thinking about ML applications, most people think of a single model with fixed inputs and outputs. Dynamic models differ in that they take into account the differing prediction difficulty of each data input. Depending on the input, a dynamic model may adapt its preprocessing steps, computational path, or model choice all together (Zhu et al., 2021). While many types of dynamic models exist, we only discuss a selection of techniques that are mentioned later on.
@@ 3.6.1. Early Exit Models
Allow for completing a prediction without the need to pass through the full model. Motivated by the fact that some data points are easier to predict than others, early exit models run intermediate feature representations through additional output layers to check if the prediction is sufficiently correct. For example, if a model is already confident about the current prediction, it finishes early rather than continuing to the next prediction layer.
@@ 3.6.2. Gated Models
Train a smaller, approximate model whose prediction decides whether or not to invoke a larger model. Similar to early exit models, some data points are easier to predict than others. A fast, smaller model gates a larger model, for example by again using the confidence of the initial prediction. At a high-level, these are systems of models that decide whether or not to spend extra compute if a prediction is still uncertain or unknown.
@@ 4.1. Study Protocol
To capture emerging practices around efficient machine learning, we conducted semi-structured interviews (Boyce and Neale, 2006; Knott et al., 2022) with 30 ML practitioners at Apple to study how they approach model compression in their own work. Our interview questions are outlined in Appendix A. We gave participants ample time to flexibly speak to their specific work and express other topics beyond our question set that they felt were important to effectively optimizing models (Knott et al., 2022). The interviews took place between March and July 2022 with each conversation lasting from 45 minutes to 1 hour. For all interviews, one researcher lead the interview questions, while another took detailed notes. Where a participant approved, we also recorded conversations to refer back to during analysis. No compensation was given, since all participants were salaried employees. At the end of the study, we briefed participants and their teams on our results. Our study protocol was approved by an internal IRB.
@@ 4.2. Participants and Recruitment
As discussed in Section 2, in-depth interview studies with ML practitioners are uniquely suited to capture experts’ tacit knowledge for the purpose of generating new, publicly-available guidance (Amershi et al., 2019a, b; Holstein et al., 2019; Sambasivan et al., 2021; Liao et al., 2020; Bhatt et al., 2020; Yang et al., 2018). Our organization has a rare concentration of efficient machine learning experts, so we first reached out to known individuals. We then used a snowball sampling strategy to reach a broader network of people involved in efficient machine learning. To balance different perspectives, we sought practitioners working on ML for different domains and tasks (e.g., vision, language, and sensing). As the study evolved, we also chose new participants to help fill-in our largest areas of uncertainty. Our participants, listed in Table 1, include ML research scientists, engineers, and managers spanning a wide breadth of application domains. A natural saturation occurred when participants began to repeat know-how we had already recorded and began to only suggest participants we had already included.
@@ 4.3. Qualitative Data Analysis
The first two authors conducted an iterative thematic analysis method to group common motivations, challenges, and best practices of model compression into categories (Gibbs, 2007). As we conducted more interviews, we continually discussed and updated these categories to reflect our new findings. Each participant’s data and transcripts were independently reviewed and manually coded by the first two authors at least twice using inductive coding (Thomas, 2003). In total, we spent 30+ hours interviewing participants and collected 23,500+ words of additional notes outside of the audio transcripts. The final categories were split into two main sections: Section 5 “Study Results” and Section 6 “Design Opportunities for Efficient ML Tools and Interfaces.” The composition of these two sections was formed by ordering major categories and hierarchically grouping within categories when relevant.
@@ 4.4. Methodological Limitations
While we found that learning from expert practitioners was tremendously valuable, any interview study has limits for generalizability. Interview studies suffer from smaller population samples than other methods like a survey. We chose to hold in-depth exploratory conversations with each participant—which a survey cannot support as well (Knott et al., 2022). On the other hand, we did not directly observe participants conducting their efficient machine learning work, which limits the specificity of our findings. It was not possible to observe all participant’s activity or model artifacts due to the sensitivity of their work. Another concern is bias from a participant’s role in the domain (Boyce and Neale, 2006), as well as the power dynamics between interviewer and interviewees (Kvale, 2006). This study was conducted solely within one organization, therefore practitioners may hold organization-specific beliefs and practices (Schein, 1990). Despite our conscious efforts to recruit across ML application domains, we noticed a skew towards vision-based applications on images, video, and 3D scene data. While interviews were conducted prior to the public rise of LLMs in late 2022 (Movva et al., 2023), many participants had NLP optimization experience, and as discussed in Section 2.3.2, the same overall approaches for efficient machine learning are currently being applied to LLMs and other foundation model modalities.
Taking these limitations together, we caution readers to not consider the advice we detail from participants universally generalizable. However it is with confidence that we share the rich pragmatic guidance these 30 experts have to offer.
@@ 5. Study Results
To answer our primary research question RQ: How should a broader audience of HCI and ML practitioners today optimize powerful models to design on-device, ML user experiences?, we first profile who is working in model compression today, as emerged from our participant pool (Section 5.1), and give a high-level state of efficient machine learning in practice (Section 5.2). We then break down our results along a typical AI/ML development process (mlw, 2023; tds, 2018; Fayyad et al., 1996; Wirth and Hipp, 2000), and use the machine learning workflow outlined by Amershi et al. (2019a) as a reference point to map onto (in summary: model and data requirements, model development and training, and model evaluation). Similar to specifying model requirements within ML (Amershi et al., 2019a), we characterize how practitioners design on-device machine learning experiences (Section 5.3) and plan model budgets (Section 5.4). Next, we discuss how compression affects the model development and training process, and offer considerations for how people can strategize applying compression to their own work (Section 5.5). Lastly, we describe how practitioners evaluate compressed models to balance accuracy versus performance—and avoid accidental compression artifacts (Section 5.6).
Throughout the results we highlight actionable takeaways in call-out boxes . Note that some of these strategies are unique to model compression for on-device ML, while other strategies add a model compression twist to already-familiar software efficiency ideas. We include both to balance domain-specific novelty with the key advice for on-device ML.
@@ 5.1. Participants and Emergent Personas
Who does on-device ML efficiency? Of the 30 people we interviewed, participants had an average of 7.1 (max 12) years experience with ML, and an average 4.1 (max 10) years experience with efficient ML. Our participants had diverse breadth of expertise across domains, detailed in Table 1. Rather than their job title, we found that participant perspectives on efficient machine learning aligned more closely with their application focus. For the purpose of understanding practitioners contributions, we define three distinct, emergent personas that best describe our participants:
  • Compression Experts (E1–E13 in Table 1): Seasoned research scientists and experienced engineers that lead on-device ML initiatives, develop new compression techniques, and consult on machine learning optimization efforts. Given the amount of experience these people have, they often manage or lead teams. Example: An ML research scientist who has a PhD in model compression and is developing novel techniques for model optimization.
  • Machine Learning Practitioners (P1–P11 in Table 1): ML engineers and data scientists that build and deploy models on-device where certain optimization budgets must be met. These people use compression as a means to an end rather than solely studying efficient machine learning. Example: An ML engineer optimizing a model’s size to shrink it to 1MB while maintaining high accuracy.
  • Tooling Engineers (T1–T6 in Table 1): Engineers and developers that focus on building frameworks, infrastructure, and tools for efficient machine learning. Example: A software engineer building and maintaining an organization-wide tool to help others compute efficiency metrics.
@@
E1 10/12 ML Manager (✓) Efficient ML research
@@
P1 6/10 ML Manager (✓) Efficient computer vision
@@
T1 5/10 ML Manager (✓) Efficient ML tooling
@@
Comparing the personas, we see clear differences between applications. For example, compression experts are heavily (and nearly exclusively) focused on research, ML practitioners work across the most diverse set of domains (e.g., vision, NLP, sensing, multi-modal models, fairness, and hardware), and tooling engineers focus on internal tools and compression algorithm implementation. Throughout the paper, we label representative quotes from participants by their personas to illustrate the main findings from the study. The persona labels offer additional context for situating a participant’s background and perspective into the larger discussion.
5.2. The State of Efficient ML in Practice
Creating an efficient machine learning model is a “large design and constrained optimization problem” [E5], where practitioners have to weigh decisions between many model performance metrics, such as model storage size, inference latency, power usage, device temperature, and model behavioral metrics such accuracy, precision, recall, and others. To many readers, this characterization may sound like a job for an automated optimization algorithm. However, a crucial finding of our results is that automation is not yet possible, do to the degree of ML, product design, and human expertise that goes into creating on-device applications. Experts emphasized there is no single solution, “silver bullet” [E12], “turnkey solution” [E5], or “golden recipe” [E7] for successfully building efficient machine learning models. Partially, this is due to the rapidly moving-target of new state-of-the-art model architectures and new ML applications that may require a custom approach.
The expertise for creating efficient ML models is currently held by a relatively small subset of ML practitioners, where “knowledge is handed-down from the few who know how to do it successfully,” [E3]. These people are referred to as artisans and “based on earlier experience, know it’s possible” [P10] to create highly efficient models. Reducing a model’s size by half, or by an order of magnitude, takes additional time and considerations [E7]. “It’s like black magic or the dark arts, but you will get better with time,” [E7]. Readers may recognize the “dark arts” or “artisan” language from earlier days of contemporary ML (Patel et al., 2008; Hill et al., 2016), before any real widespread knowledge base was established. We hope these study results support growing such a knowledge-base for efficient machine learning, and it’s role within a holistic ML process:
[Efficient ML] means a lot of things to different people. Specific techniques on models weights help reduce size, but to get an efficient model comes from more careful design of the loss function, the system, which parts should and should not be modeled with ML.” — E5
5.3. The Metrics of Efficient, On-Device Models
@@
All metrics: model tested during high device resource load Resources are shared. A user’s device may be running many things simultaneously, including multiple ML models. A single model that demands too many resources will slow everything down.
@@
As an umbrella term, efficiency encompasses many metrics. Common metrics that practitioners linked to user experience are summarized in Table 2. Some may be familiar with metrics commonly considered in ML workflows, such as latency and model storage size. Other metrics may come as more foreign to an ML engineer accustomed to deploying their model server-side, e.g., device temperature. When a user is holding and carrying around a device all day, metrics like heat and battery life impact are crucial.